OpenStudy (anonymous):

Use your knowledge of the derivative to compute the value of the limit given below: \lim_{h \rightarrow 0} \frac{\sqrt[7]{(x + h)^{8}} - \sqrt[7]{x^{8}}}{h} .

OpenStudy (anonymous):

I just hope I typed this correctly. I copied and pasted from a problem that was similar to mine, I just changed out the numbers I needed. It's for the seventh root of x^8

OpenStudy (anonymous):

\[ \lim_{h \rightarrow 0} \frac{\sqrt[7]{(x + h)^{8}} - \sqrt[7]{x^{8}}}{h} \]

OpenStudy (anonymous):

Yea, that's it.

OpenStudy (anonymous):

Remember that \[ f'(x) = \lim_{h \rightarrow 0} \frac{f(x + h) - f(x)}{h} \]And in this case: \[ f(x) = \sqrt[7]{x^{8}} \]

OpenStudy (anonymous):

I understand that, but I'm unsure how to solve it. If I remember correctly I have to multiply by it's conjugate (I hope that's the correct word), but I keep doing it incorrectly.

OpenStudy (anonymous):

You just have to take the derivative....\[ \Large f(x) = \sqrt[7]{x^8} = x ^{\frac{8}{7}} \]Where is \(f'(x)\)?

OpenStudy (anonymous):

got it, thank you